# Universal Dense-Matter Trace Anomaly Inferred from Collective Flow in Heavy-Ion Collisions and Global Properties of Neutron Stars

### v1

@@ Abstract
where $\bar{\varepsilon}\equiv\varepsilon/\varepsilon_{0}$ and $\varepsilon_{0}\simeq 150~\mathrm{MeV/fm}^{3}$ denotes the energy density at the saturation density $\rho_{0}$. The first term on the right-hand side represents the derivative contribution, while the remaining terms constitute the non-derivative part. Figure 4 compares the resulting $c_{s}^{2}(\varepsilon)$ inferred from heavy-ion collisions with constraints from neutron star observations. Despite comparatively larger uncertainties arising from the derivative nature of Eq. (7) and the finite sampling of beam energies in the GSI data, the inferred $c_{s}^{2}(\varepsilon)$ values are consistent with the neutron star constraints. Overall, the mean values of $c_{s}^{2}(\varepsilon)$ from heavy-ion collisions cluster closely around the best-fit sound-speed profile obtained independently from neutron star observations using machine-learning techniques [3].
Taken together, terrestrial heavy-ion collisions and neutron star observations reveal a universal macroscopic response of dense baryonic matter, characterized by a nearly model-independent trace anomaly and a consistent sound-speed profile over the overlapping range of energy densities probed by both systems.
@@
Conclusions.— In summary, we have performed the first Bayesian extraction of the dense-matter trace anomaly from collective flow observables in intermediate-energy heavy-ion collisions. Although these collisions generate hot, dynamically evolving matter, our transport simulations successfully decouple the cold-matter EOS through the nucleon mean-field potential from separate thermal effects handled via collision terms. This approach allows collective flow data to directly constrain macroscopic properties of the cold dense-matter EOS.
The trace anomaly inferred from these laboratory experiments shows remarkable quantitative agreement with independent constraints derived from neutron star observations, including mass-radius relations and tidal deformabilities. This consistency is particularly significant given the vastly different data sets, priors, and inference frameworks employed.
Such nontrivial agreement demonstrates that collective flow in heavy-ion collisions and global neutron star observables probe the same coarse-grained macroscopic properties of dense matter, despite their vastly different dynamical and thermodynamic conditions. This result establishes collective flow as a laboratory-based, model-agnostic probe of dense matter that is directly complementary to neutron star observations.
Looking forward, upcoming precision measurements at next-generation heavy-ion facilities, combined with more refined neutron star data from advanced X-ray timing and gravitational-wave detectors, will further sharpen our understanding. These advancements will provide critical tests of the universality of the trace anomaly and map the dense-matter EOS across even broader ranges of energy and density.
@@
DATA AVAILABILITY: The data that support the findings of this article will be openly available [51].

### v2

@@ Abstract
Figure 3 compares the resulting $c_{s}^{2}(\rho)$ inferred from heavy-ion collisions with constraints from neutron star observations. Despite comparatively larger uncertainties arising from the derivative nature of Eq. (4) and the finite sampling of beam energies in the GSI data, the inferred $c_{s}^{2}(\rho)$ values are consistent with the neutron star constraints. Overall, the mean values of $c_{s}^{2}(\rho)$ from heavy-ion collisions cluster closely around the best-fit sound-speed profile obtained independently from neutron star observations using machine-learning techniques [3].
Taken together, neutron star observables and heavy-ion collective flow provide independent probes of dense matter that are primarily sensitive to the macroscopic EOS. Their convergence indicates a consistent macroscopic EOS over the explored density range, establishing the trace anomaly as a bridge observable linking terrestrial experiments and astrophysical observations.
@@
Conclusions.– In summary, we have demonstrated that collective flow observables in intermediate-energy heavy-ion collisions enable the first Bayesian extraction of the cold dense-matter trace anomaly, $\Delta(\varepsilon)$, and the corresponding EOS parameter $w\equiv P/\varepsilon$. Although these reactions create hot and dynamically evolving systems, our transport simulations enable a practical decoupling of the cold-matter EOS—encoded in the nucleon mean-field potential—from thermal effects treated through collision terms. Within this framework, collective flow data constrain the macroscopic stiffness of the cold dense-matter EOS. Because collective flow develops continuously during the reaction and is generated by pressure gradients integrated over the space–time evolution of the system, the inferred trace anomaly at a given beam energy represents an effective average of the EOS stiffness over densities up to the maximum compression reached in the collision.
The trace anomaly inferred from these laboratory experiments shows remarkable quantitative consistency with independent constraints derived from global neutron star observables. This agreement is nontrivial, given the very different data sets, priors, and inference frameworks. It indicates that collective flow and neutron star global observables probe the same macroscopic properties of cold dense matter despite their very different dynamical conditions. Our results demonstrate that collective flow provides a robust laboratory probe of the macroscopic EOS that is complementary to neutron star measurements. By exploiting the beam-energy dependence of collective flow, which corresponds to different maximum compression densities reached in the collision, the trace anomaly can be mapped over a range of energy densities and used to reconstruct the density dependence of the EOS stiffness. Future high-precision experiments at next-generation heavy-ion facilities, together with improved constraints from advanced X-ray timing and gravitational-wave observations of neutron stars, will allow stringent tests of the dense-matter trace anomaly and further refine constraints on the EOS over a broad range of densities. These developments will further establish the trace anomaly of cold dense matter—constrained by collective flow—as a bridge observable linking terrestrial heavy-ion experiments and neutron star observations.
@@
DATA AVAILABILITY: The data supporting the findings of this work, including processed data and posterior samples, will be made openly available [75].